OpenStudy (anonymous):

x^5y^3z^3 ---------- xy^-3z^-1 Which of the following is the simplest form of this expression? Select one of the options below as your answer: A. x^4z^2 B. x^4y^6z^4 C. x^5y^-1z-3 D. x^6y^6z^4

OpenStudy (anonymous):

B

OpenStudy (campbell_st):

when dividing subtract the powers of lik pronumerals \[x^{5 -1} = x^4\] \[y^{3 - -3} y^6\] \[z^{3 - -1} = z^4\] the solution is \[x^4y^6z^4\]